OpenStudy (anonymous):

Salut, J'ai un exercice: Montrer par récurrence que x^n+1/x^n=2cos(nx) n appartient N/{0}. Sachant que x+1/x=2cosx

OpenStudy (anonymous):

hello :)

OpenStudy (anonymous):

do you mean : show that\[x^n+\frac{1}{x^n}=2 \cos (n \theta)\]if\[x+\frac{1}{x}=2 \cos \theta\]?

OpenStudy (anonymous):

oui

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

Who can help me ?

OpenStudy (inkyvoyd):

is n a positive integer? and will proof by mathematical induction do?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let suppose statement holds for \(n\) \[x^n+\frac{1}{x^n}=2 \cos (n \theta)\]so using mathematical induction we must prove it holds also for \(n+1\) or in mathematical words\[x^{n+1}+\frac{1}{x^{n+1}}=2 \cos ((n+1) \theta)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

very well :)

OpenStudy (anonymous):

now i give u a hint

OpenStudy (anonymous):

are u supposed to use mathematical induction for sure?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you continue the calculation

OpenStudy (anonymous):

well yes

OpenStudy (anonymous):

using 2 statement that we know they are right\[\left(x+\frac{1}{x} \right) \left(x^n+\frac{1}{x^n} \right)=4 \cos \theta \cos (n\theta)\]am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

continue pleaze

OpenStudy (anonymous):

ok distribute the LHS \[x^{n+1}+\frac{1}{x^{n+1}}+x^{n-1}+\frac{1}{x^{n-1}}=4 \cos \theta \cos (n\theta)\]note that we use strong induction so\[x^{n-1}+\frac{1}{x^{n-1}}=2 \cos((n-1)\theta)\]put this in the latest equation u will get\[x^{n+1}+\frac{1}{x^{n+1}}+2 \cos((n-1)\theta)=4 \cos \theta \cos (n\theta)\]\[x^{n+1}+\frac{1}{x^{n+1}}=4 \cos \theta \cos (n\theta)-2 \cos((n-1)\theta)\]now you should prove RHS is equal to \(2 \cos((n+1)\theta)\) for completing the proof

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

can you continue pleaze ?

OpenStudy (anonymous):

using formula \[\cos A \cos B=\frac{1}{2}(\cos(A+B)+\cos(A-B))\]we will have\[4 \cos \theta \cos (n\theta)-2 \cos((n-1)\theta)=2(\cos((n+1)\theta)+\cos((n-1)\theta)-2 \cos((n-1)\theta)\]\[4 \cos \theta \cos (n\theta)-2 \cos((n-1)\theta)=2 \cos((n+1)\theta)\]

OpenStudy (anonymous):

thanks !!

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

actually mathematical induction is not what i interested in for this problem

OpenStudy (anonymous):

i give you a solution for fun :)

OpenStudy (anonymous):

do you know complex numbers?

OpenStudy (anonymous):

discuss the values ​​of these two equations: 1) x ²-2xsina-cos ² a = 0 2) ² x ²-2x + cos a = 0

OpenStudy (anonymous):

pleaze this exercice

OpenStudy (anonymous):

post it in open questions, i gotta go :)